# A Unified Instance Segmentation Framework for Completely Occluded Objects and Dense Objects in Robot Vision Measurement

### v2

@@ IV-C Completely Occluded Object Instance Segmentation
Method
AP
AP50
AP75
FPS
AmodalMask [15]
19.49
35.18
24.36
16.1
ASN [5]
40.88
61.24
45.93
18.6
ORCNN [6]
37.60
57.78
40.16
22.2
BCNet [4]
45.30
65.13
51.79
7.4
@@
50.09
69.30
56.42
2.7
SAM [13]
26.15
42.74
30.28
2.8
Fine-tuned SAM
68.66
82.23
71.38
2.8
CFNet*
75.13 89.21 77.98
2.7
CFNet
73.46
86.48
76.20
20.4
@@ IV-D Dense Object Instance Segmentation
Method
AP
AP50
AP75
FPS
OIS [10]
27.41
47.85
30.71
18.4
RB Mask R-CNN [9]
29.07
50.96
32.64
16.3
ISOP [12]
34.73
59.16
36.84
19.6
@@
37.81
61.95
39.45
2.7
SAM [13]
14.37
41.88
18.52
2.8
Fine-tuned SAM
20.26
45.30
24.59
2.8
CFNet*
45.10 66.45 47.92
2.7
CFNet
43.76
64.29
46.13
20.4
@@ IV-E2 Knowledge Distillation
We first compare the performance of CFNet learning from the teacher model and GT labels, and test the impact of different label smoothing methods. “T” represents learning from the teacher model. The smoothing factor in ULS is 0.1. Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0. The weight factor of the teacher model and GT labels is 0.1. “${}^{-}$” represents distillation only for the mask decoder.
@@
$\delta=0.5$
$\delta=1.0$
$\delta=0.5$
$\delta=1.0$
$k=3$
64.24
64.24
64.20
64.17
$k=5$
64.22
64.29
64.18
64.21
$k=7$
64.19
64.25
64.15
64.14
@@ IV-E3 Oriented Bounding Box Detection
Mask
Oriented
Oriented
S2ANet [27]
R-CNN [16]
R-CNN [14]
Reppoints [26]
42.95
64.29
64.37
64.06

### v3

# An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes

@@ IV-C Completely Occluded Object Instance Segmentation
Method AP AP50 AP75 FPS
AmodalMask [15] 19.49 35.18 24.36 16.1
ASN [5] 40.88 61.24 45.93 18.6
ORCNN [6] 37.60 57.78 40.16 22.2
BCNet [4] 45.30 65.13 51.79 7.4
@@
SAM [13] 26.15 42.74 30.28 2.8
Fine-tuned SAM 68.66 82.23 71.38 2.8
CFNet* 75.13 89.21 77.98 2.7
CFNet 73.46 86.48 76.20 20.4
@@ IV-D Dense Object Instance Segmentation
Method AP AP50 AP75 FPS
OIS [10] 27.41 47.85 30.71 18.4
RB Mask R-CNN [9] 29.07 50.96 32.64 16.3
ISOP [12] 34.73 59.16 36.84 19.6
@@
SAM [13] 14.37 41.88 18.52 2.8
Fine-tuned SAM 20.26 45.30 24.59 2.8
CFNet* 45.10 66.45 47.92 2.7
CFNet 43.76 64.29 46.13 20.4
@@ IV-E2 Knowledge Distillation
We first compare the performance of CFNet learning from the teacher model and GT labels, and test the impact of different label smoothing methods. “T” represents learning from the teacher model. The smoothing factor in ULS is 0.1. Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0. The weight factor of the teacher model and GT labels is 0.1. “GS-” represents distillation only for the mask decoder.
@@
$\delta=0.5$ $\delta=1.0$ $\delta=0.5$ $\delta=1.0$
$k=3$ 64.24 64.24 64.20 64.17
$k=5$ 64.22 64.29 64.18 64.21
$k=7$ 64.19 64.25 64.15 64.14
@@ IV-E3 Oriented Bounding Box Detection
Mask Oriented Oriented S2ANet [27]
R-CNN [16] R-CNN [14] Reppoints [26]
42.95 64.29 64.37 64.06